OpenStudy (anonymous):

which of the following is a counter example for the given statement 2^n+1 is divisible by 3 a=1 b=3 c=2 d=5

OpenStudy (anonymous):

Let's see! 2^1+1=3 (divisible by 3) 2^3+1=9 (divisible by 3) 2^2+1=? 2^5+1=? Give it a try! Tell us what do you get please!

OpenStudy (anonymous):

okay thank you !

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

@alex1601 Could you please tell us what did you get? We are here to help you!

OpenStudy (anonymous):

ok so it wasent a so im thinking its b

OpenStudy (anonymous):

b isn't because of this: 2^3+1=8+1=9 (divisible by 3) What about c? 2^2+1=?

OpenStudy (anonymous):

okay ill try that one and ill tell you in a bit if it was right

OpenStudy (anonymous):

Great! Take your time.

OpenStudy (anonymous):

rafabc02 yes the answer was c thank you !

OpenStudy (anonymous):

You are welcome! Thanks for the "best response"! Have a great night!

OpenStudy (anonymous):

welcome! ..you too !